# A Survey on Efficient Federated Learning Methods for Foundation Model Training

### v1

@@ Abstract
Federated Learning (FL) has become an established technique to facilitate privacy-preserving collaborative training. However, new approaches to FL often discuss their contributions involving small deep-learning models only. With the tremendous success of transformer models, the following question arises: What is necessary to operationalize foundation models in an FL application? Knowing that computation and communication often take up similar amounts of time in FL, we introduce a novel taxonomy focused on computational and communication efficiency methods in FL applications. This said, these methods aim to optimize the training time and reduce communication between clients and the server. We also look at the current state of widely used FL frameworks and discuss future research potentials based on existing approaches in FL research and beyond.
{textblock*}
5cm(1.52cm,26cm) Preprint. Under Review.
@@ 1 Introduction
Transformers, Foundation Models (FMs), and Large Langauge Models have conquered the deep learning world with unprecedented speed, enabling generative artificial intelligence for a broad audience. The most prominent example is OpenAI’s ChatGPT OpenAI (2023). Yet, these models require a carefully curated training dataset with several billion, or even trillion, samples. This limits the number of organizations capable of pre-training FMs to a handful of enterprises Zhou et al. (2023). Yet, fine-tuning methods for FMs have been introduced to leverage pre-trained FMs for their specific purposes (downstream tasks), which only require a fraction of the amount of data necessary for pre-training Kalyan et al. (2021). However, in practice, access to data is oftentimes challenging due to privacy regulations or environmental circumstances, such as poor connectivity between cloud and network edge devices Osia et al. (2020); Shokri and Shmatikov (2015).
To enable learning from massively distributed data and poorly connected clients with $<1\mathrm{GBit/s}$ bandwidth, McMahan et al. [2017] introduced Federated Learning (FL), a privacy-preserving machine learning technique that does not require the transfer of input data to a central server. However, this often comes at the cost of having non-identically and non-independently distributed (non-IID) data, variable client computational resources, and unstable communication, up to a point where the time spent on computation is as high as the time spent on communication Yousefpour et al. (2023). For FL applications, it is desirable to maximize computation and minimize communication. FMs, due to their large number of parameters Bommasani et al. (2021), already tend to take significantly longer to fine-tune than the models that are widely used in FL research to date He et al. (2020). As such, we see a natural focus on computation, which, in principle, can benefit us during training Ryabinin et al. (2023); Woisetschläger et al. (2023); Beutel et al. (2020). That said, it is more likely to use FL systems to fine-tune pre-trained FMs by, e.g., applying parameter-efficient fine-tuning techniques (PEFT) rather than pre-training. Pre-training generally requires full control over the training data and high data quality, which is hard to ensure in federated settings Longpre et al. (2023). PEFT often reduces the number of trainable parameters below 1% of all model parameters. As such, PEFT also helps reduce communication as we only need to transmit trainable parameters (see Section 3.2). This contrasts previously existing literature on efficiency methods for FL. There, computation and communication are perceived as separate topics, and experiments were carried out with small, full models.
@@
Our survey is the first to study advances in computational and communication-efficient methods in FL applications with a special focus on FM training. Our work contains three distinct contributions:
  • A novel taxonomy on FMs in FL systems based on the core challenges in computation and communication. Our taxonomy focuses on computational and communication advancements that benefit FM training and fine-tuning.
  • In-depth discussion of current state computational and communication efficient methods for FL and FMs. FMs are large transformer models, often having several billions of parameters, making them expensive to communicate, train, or fine-tune. We discuss methods to facilitate efficient on-device training and fine-tuning. We also look into modern compression techniques for communicating large models.
  • A thorough overview of future research directions. We highlight future directions for research on computational and communication efficiency with regard to the harmonized evaluation of new methods. We further point out the increased complexity of hyperparameter optimization with PEFT and briefly touch upon the differences in privacy for full models vs. partial model updates for FMs.
@@ 2 Taxonomy
Our taxonomy introduces a novel perspective focused on the current developments in the field of efficient computational and communication methods for FL applications intended for training and fine-tuning large transformer models. While communication efficiency has been studied extensively in FL, computational advancements in conjunction with large models are currently emerging. Our taxonomy is visualized in Figure 1.
@@ 2.1 Basics of Federated Learning
Coordinated by a server, the overall goal of FL is to collaboratively and iteratively train a DL model across a set of clients $N$ and minimize the global loss, where $f$ denotes an objective function with parameters $w$,
$\min_{w}\,f(w)\mathrel{\mathop{:}}=\frac{1}{|N|}\sum_{n=1}^{|N|}f_{n}(w)\mathrm{.}$ (1)
@@ 2.2 Taxonomy Explanation
Our taxonomy is centered on two major challenges in FL: the computational and communication efficiency layers. With the emergence of FMs in FL applications, we see computational needs, as well as communication requirements, increase exponentially since models for real-world applications had a few 100K to 10M parameters in the past Wang et al. (2023); Huba et al. (2021). Now, models often reach billion-scale parameter sizes, i.e., a growth factor of more than 100×100$\times$100 × Zhang et al. (2023a).
Computational Efficiency. For smaller models, computation times on FL clients are seconds, while for large models, training can take several hours or even days. There are various techniques to lower computation times for FMs by reducing the number of trainable parameters.
Specifically, we discuss (I) full model training in FL systems as the longest-standing paradigm in FL (Section 3.1), (II) parameter-efficient fine-tuning as the most common technique for transformer fine-tuning (Section 3.2), (III) prompt tuning as a highly efficient method without training any model parameters (Section 3.3), and (IV) instruction tuning as a method to train a single model on heterogeneous downstream tasks (Section 3.4). For each technique, we outline its benefits and drawbacks, especially with regard to data heterogeneity.
@@
Communication Efficiency. While computationally efficient methods for FM training in FL applications already reduce the number of parameters and data points to communicate, there is still a significant amount of data to be communicated between clients and servers. For instance, parameter-efficient fine-tuning methods only require 1–2% of parameters to be trainable. This still amounts to up to 14M parameters when working with Alpaca-7B Zhang et al. (2023a), larger than the majority of models currently being discussed in FL research He et al. (2020); Beutel et al. (2020).
@@
We discuss communication efficiency methods along two major categories. (I) Model pruning is a method to communicate parts of a model between clients and the server, which resemble the most important parameters for a client (Section 4.1). (II) Full model compression is divided into three sub-categories: First, quantization is a method to decrease the numeric precision of model parameters. Second, sparsification is used as a way to zero out less important model parameters. Third, low-rank compression transforms high dimensional parameter matrices in client updates into significantly lower dimensional representations (Section 4.2).
Table 1: Computational Efficiency Methods for FL Systems and FM. Foundation Models (FMs) generally provide a strong performance across a variety of domains that are well explored in centralized learning Dosovitskiy et al. (2020); OpenAI (2023); Yang et al. (2023); Liu et al. (2023).
Training Underlying FL Max. Model Domain
@@
Paper Regime Aggregation Strategy Params CV NLP Audio Graph
Centralized Learning FMT, PEFT, PT, IT $>100\mathrm{B}$
FedBERT Tian et al. (2022) FMT Weighted Average 117M
FedCLIP Lu et al. (2023) PEFT Weighted Average 85M
FedPEFT Sun et al. (2023) PEFT Weighted Average 85M
FedPETuning Zhang et al. (2023b) PEFT Weighted Average 125M
SLoRA Babakniya et al. (2023) PEFT Weighted Average 67M
FedPrompt Zhao et al. (2022) PT Weighted Average 223M
FedIT Zhang et al. (2023a) IT Weighted Average 7B
3 Federated Learning Methods for FMs
@@ 3.1 Full Model Training
Generally, we refer to full model training when all parameters of a neural network are being trained. This is particularly useful when permanently fitting a model to a specific use case. In this approach, we locally train a model on all clients with the objective of minimizing $l$.
One of the first foundation model architectures is the BERT transformer. Tian et al. [2022] discuss federated pre-training of BERT-family models with up to 117M parameters for masked language modeling (MLM) and next-sentence prediction tasks. In MLM, the loss is defined over the sum of probabilities $P$ of predicted tokens $\hat{x}$ over the representation function of a masked sentence $g\big{(}\frac{x}{M(x)}\big{)}$,
@@
MLM is the prevailing technique for self-supervised learning as it accelerates training on massive language bodies Raffel et al. (2019). Especially in federated environments, this enables one to learn from user data without the necessity of an expensive or manual labeling process.
Results show the effectiveness of federated pre-training and provide a net benefit over training on local datasets only Babakniya et al. (2023); He et al. (2020). However, the perplexity for large models (e.g., GPT-2) trained in a federated fashion is four orders of magnitude worse than centralized training, which is a stark indicator of low model quality Tian et al. (2022). However, for use cases involving sensitive data and strict privacy regulation, full model pre-training opens up the opportunity to create foundation models based on federated data.
Though full model training offers the highest degree of customization to a model, the technique requires substantial computational resources and communicating the entire model back and forth between clients and the server. In a federated setting, this can be very expensive and only feasible for a small number of participants. Therefore, more efficient methods are required.
@@ 3.2 Parameter Efficient Fine-Tuning
Generally, PEFT is used for large models that were trained on a large data basis already, provide good performance across a great variety of tasks, and require fine-tuning on a dataset for a specific downstream task to deliver excellent performance. This is especially effective since the data required for fine-tuning is orders of magnitude smaller than for pre-training. Also, since we use a pre-trained model and freeze the model parameters, the sensitivity towards non-IID data in fine-tuning is significantly lower than in pre-training. This renders PEFT techniques well-suited for FL applications and explains why this is a frequently discussed topic. Aside from the positive effects with regard to non-IID data, PEFT also reduces the computational costs for fitting a model to a specific downstream task. To achieve this, there are two methods as follows.
Sparse fine-tuning of pre-trained model parameters. As communication is a key concern in FL, it has become a priority to reduce the number of parameters to communicate between client and server. One of the most used approaches to achieve this is BitFit Zaken et al. (2021). The technique freezes almost all model parameters $w_{t}$ and only trains the bias term $b$ and the final classification layer $w^{\mathrm{final}}_{t}$ over the input features $a_{t}$, where the next-layer input features $a_{t+1}=a_{t}\cdot w_{t}+b$.
With this technique, it is only required to communicate $b$ and $w^{\mathrm{final}}_{t}$. Communication efforts are reduced by $\geq 99\%$, i.e., instead of communicating several 100M parameters, only a few 100K parameters are communicated.
@@
Adapter-based fine-tuning of additionally added parameters. When aiming to maintain a pre-trained model while introducing task-specific knowledge, adapter-based fine-tuning techniques provide strong performance and on-par efficiency as compared to sparse fine-tuning techniques. Here, we introduce two adapter layers in each transformer block of a foundation model. The adapter layer $a_{t+1}$ is calculated based on a downstream projection $w_{t}^{\mathrm{down}}$ of input feature $a_{t}$ into a lower-dimensional space $w_{t}^{\mathrm{down}}\in\mathcal{R}^{d\times r}$ followed by an upstream projection $w_{t}^{\mathrm{up}}\in\mathcal{R}^{r\times u}$, resulting in
$a_{t+1}=w_{t}^{\mathrm{up}}\cdot h(w_{t}^{\mathrm{down}}\cdot a_{t})\mathrm{.}$ (4)
With this, we can fine-tune an FM over much fewer dimensions than when fully fine-tuning a model. This saves both computational resources and communication on a client.
An improvement with regard to computational and communication efficiency over additional adapters is low-rank adapters (LoRA) Hu et al. (2021). The technique uses a lower dimensional representation $\mathbf{A}\in\mathcal{R}^{r\times u}$ and $\mathbf{B}\in\mathcal{R}^{d\times r}$ for weight updates $w_{t+1}=w_{t}+\Delta w=w_{t}+\mathbf{B}\mathbf{A})$. $r\ll\min(d,u)$ casts the weight update matrices into a much lower dimensionality than in the original transformer module without the necessity of adding additional adapters, i.e., LoRA builds an adapter for existing parameters. However, as $\mathbf{A}$ is initialized randomly to a Gaussian distribution and $\mathbf{B}$ as a zero matrix, this works well for centralized settings with IID data Hu et al. (2021). For FL settings, this initialization method bears the risk of slowing down the fine-tuning process over non-IID data.
@@
Zhang et al. [2023b] provide a systematic benchmark study on adapter-based fine-tuning methods in privacy-preserving FL systems. Their results show that fine-tuning with additional adapters and LoRA both yield the same benchmark results regarding model accuracy. However, LoRA requires 3×3$\times$3 × less communication than additional adapters.
However, both adapter-based methods yield a worse accuracy than full fine-tuning. With SLoRA, Babakniya et al. [2023] propose a method to optimize LoRA for non-IID FL settings. Their approach entails parameterizing the weight update based on $r$,
@@
As $\beta\in r$, scaling the fraction helps control the weight update impact of a single client. Subsequently, this can be used to control the negative effects caused by non-IID data. To practically achieve this, Babakniya et al. [2023] make use of a two-stage process: First, they used singular vector decomposition on $\mathbf{A}$ and $\mathbf{B}$ to obtain a common initialization point for LoRA across all clients in an FL system. Second, the training is facilitated with the commonly initialized low-rank representations. Their approach achieves on-par performance with full model fine-tuning. However, they require a warmup time of approximately 100 FL rounds for stage 1, which can be very expensive in FL settings as clients are often unavailable consecutively for such a long time McMahan et al. (2017).

### v3

@@ Abstract
Federated Learning (FL) has become an established technique to facilitate privacy-preserving collaborative training across a multitude of clients. However, new approaches to FL often discuss their contributions involving small deep-learning models only and focus on training full models on clients. In the wake of Foundation Models (FM), the reality is different for many deep learning applications. Typically, FMs have already been pre-trained across a wide variety of tasks and can be fine-tuned to specific downstream tasks over significantly smaller datasets than required for full model training. However, access to such datasets is often challenging. By its design, FL can help to open data silos. With this survey, we introduce a novel taxonomy focused on computational and communication efficiency, the vital elements to make use of FMs in FL systems. We discuss the benefits and drawbacks of parameter-efficient fine-tuning (PEFT) for FL applications, elaborate on the readiness of FL frameworks to work with FMs and provide future research opportunities on how to evaluate generative models in FL as well as the interplay of privacy and PEFT.
[…]
Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
@@ 1 Introduction
Foundation Models (FMs) Bommasani et al. (2021) have conquered the deep learning world with unprecedented speed, enabling generative artificial intelligence for a broad audience. As FMs have been pre-trained on an extensive data basis and can be used in multi-modal applications, they perform well over a wide range of tasks. To specialize these models on a downstream task, we use fine-tuning that can either be done over the full model or with parameter-efficient fine-tuning techniques (PEFT) Hu et al. (2021); Lester et al. (2021); Zaken et al. (2021). A major advantage is that fine-tuning requires orders of magnitude smaller datasets than pre-training but benefits from access to a variety of samples pertaining to a task.
Access to a breadth of data has always been challenging in deep learning, as data owners are typically reluctant to share their data with service providers. To tackle the data access challenge, McMahan et al. [2017] introduced Federated Learning (FL). FL enables privacy-preserving machine learning over decentralized data without the necessity of sharing input data and does not require high bandwidth client connections. Rather, a set of clients collectively train a model and send their local model updates to a server that subsequently aggregates the updates to a global model. In FL applications that involve small models with less than 1 million parameters, we may spend as much time on communication as we spend on computation Yousefpour et al. (2023). However, it is desirable to design systems in such a way that computation takes up the majority of time. Luckily, the larger the models become, the time spent on training is larger than on communication Ryabinin et al. (2023); Woisetschläger et al. (2023); Beutel et al. (2020). As such, scaling the model size in FL systems can be desirable, and this introduces beneficial properties that can aid us in training large models with several 100 million parameters and beyond. These properties render FL the perfect choice for fine-tuning FMs for downstream tasks. FL can provide access to a significantly larger and more diverse data basis while benefiting from the increased time spent on the computation of FMs over small models. Yet, the costs of transmitting model updates remain significant even with the increased computational load of FMs Yousefpour et al. (2023), making it a priority to jointly optimize the training and communication efficiency.
@@
Our survey is the first to study advances in computational and communication-efficient methods for FM training in FL applications. Our work contains three distinct contributions:
  • A novel taxonomy on FL methods for FM training focused on the core challenges in computation and communication. We discover a gap between FL methods to increase computational efficiency and techniques to improve communication efficiency. While we see research on computational efficiency for FM training and fine-tuning in FL applications, communication efficiency methods are predominantly tailored to full-model training. Our taxonomy aims to identify synergies between FL methods for FMs and efficient communication methods.
  • Holistic evaluation of existing FL computational efficiency methods for FMs and communication efficiency techniques. We study how existing techniques can help drive the adoption of FMs in FL applications and what needs to be done to render FL frameworks ready for large models.
  • We thoroughly discuss future research directions. We highlight future directions for research on computational and communication efficiency as these domains grow closer together. Also, we show what is necessary to make FMs in FL applications a reality, especially with regard to generative tasks and privacy considerations.
@@ 2 Taxonomy
Our taxonomy introduces a novel perspective focused on the current developments in the field of efficient computational and communication methods for FL applications intended for training and fine-tuning FMs. While communication efficiency has been studied extensively in FL, computational advancements in conjunction with large models are currently emerging. Our taxonomy is visualized in Figure 1.
@@ 2.1 Basics of Federated Learning
Usually coordinated by a server, the overall goal of FL is to collaboratively and iteratively train a DL model across a set of clients $N$ and minimize the global loss, where $f$ denotes an objective function with parameters $w$,
$\min_{w}\,f(w):=\frac{1}{|N|}\sum_{n=1}^{|N|}f_{n}(w)\mathrm{.}$ (1)
@@ 2.2 Taxonomy Explanation
Our taxonomy is centered on two major challenges in FL: the computational and communication efficiency levers. While these challenges have been studied in FL extensively in the past Zhang et al. (2021), existing approaches predominantly focus on small models with $<10\mathrm{M}$ parameters. With the emergence of FMs as the backbone for multi-modal applications, we need to combine computational and communication efficiency in FL as these FMs typically come with $>1~{}\mathrm{billion}$ parameters Zhang et al. (2023a); a growth factor of 100×100$\times$100 × This introduces additional computational load for FL clients, while their resources are often scarce already Beutel et al. (2020). At the same time, communication loads are also growing since many parameters need to be transmitted.
Computational Efficiency.
We discuss computational efficiency levers along four major categories. Full model training is used to train large transformer models from the very beginning (Section 3.1). Parameter-efficient fine-tuning techniques can be utilized to improve a pre-trained FM for a specific downstream task (Section 3.2). Prompt tuning enables performance improvements of an FM without training the model itself but by designing textual prompts that we prepend to an input (Section 3.3). Instruction tuning enables fine-grained control over the model training process and allows for a high degree of model specialization for certain downstream tasks (Section 3.4).
Communication Efficiency.
@@
While computationally efficient methods for FM training in FL applications can already reduce the number of parameters to communicate, there is still a significant amount of data to be communicated between clients and servers. For instance, parameter-efficient fine-tuning methods only require 1–2% of parameters to be trainable. This still amounts to up to 14M parameters when working with Alpaca-7B Zhang et al. (2023a), larger than the majority of models (with $<1\mathrm{M}$ trainable parameters) currently being discussed in FL research He et al. (2020); Beutel et al. (2020).
@@
We discuss communication efficiency methods along two major categories. (I) Model pruning is a method to communicate parts of a model between clients and the server, which resemble the most important parameters for a client (Section 4.1). (II) Full model compression is divided into three sub-categories: First, quantization is a method to decrease the numeric precision of model parameters. Second, sparsification is used as a way to zero out less important model parameters. Third, gradient projection transforms high dimensional parameter matrices in client updates into scalar vectors for communication (Section 4.2).
Training Underlying FL Max. Model Trainable Comms Savings Domain
@@
Paper Regime Aggregation Strategy Params Parameters compared to FMT CV NLP Audio
Centralized Learning FMT, PEFT, PT, IT $\geq 7\mathrm{B}$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919 Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919 Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FedBERT Tian et al. (2022) FMT Weighted Average 117M 110M 0% Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FedCLIP Lu et al. (2023) PEFT Weighted Average 85M 530K $>99\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FedPEFT Sun et al. (2023) PEFT Weighted Average 85M 230K $>99\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FedPETuning Zhang et al. (2023b) PEFT Weighted Average 125M 1.25M $99\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
SLoRA Babakniya et al. (2023) PEFT Weighted Average 67M 140K $>99\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FedDPA Yang et al. (2024) PEFT Weighted Average 7B N/A $>99\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FedPrompt Zhao et al. (2022) PT Weighted Average 223M $>99\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FedIT Zhang et al. (2023a) IT Weighted Average 7B $>99\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
Table 1: Computational Efficiency Methods for FL Systems and FM. FMs are generally multi-modal and provide a strong performance across a variety of domains that are well explored in centralized learning but not in FL Dosovitskiy et al. (2020); OpenAI (2023); Yang et al. (2023).
3 Computational Efficiency
@@ 3.1 Full Model Training
Generally, full model training is referred to when training all parameters of a neural network. In this approach, we locally train a model on all clients with the objective of minimizing the loss $l$.
The BERT model is one of the first models to use the transformer layer architecture – the building blocks of FMs – to achieve state-of-the-art performance at the time of its release. Tian et al. [2022] discuss federated pre-training of BERT-family models with up to 117M parameters by applying masked language modeling (MLM). In MLM, the loss is defined over the sum of probabilities $P$ of predicted tokens $\hat{x}$ over the representation function of a masked sentence $g\big{(}\frac{x}{M(x)}\big{)}$,
@@
In FL applications, supervised learning can be challenging as we cannot ensure a proper data labeling process for supervised learning because we usually cannot access client data. Here, MLM can be beneficial since it is a self-supervised learning technique that masks parts of a sentence. Those masked parts will be used as the prediction target to automatically create an input and target sample. Federated pre-training provides a net benefit over training on local datasets Ding et al. (2023); Babakniya et al. (2023); He et al. (2020). However, the perplexity, an indicator for quality in large models, yields four orders of magnitude worse results for large models (e.g., GPT-2) trained in a federated fashion than centralized training, which is a stark indicator of low model quality Tian et al. (2022); Radford et al. (2019). Nonetheless, for use cases involving sensitive data and strict privacy regulation, full model pre-training allows the creation of foundation models based on federated data.
While FL pre-training has shown some promise, it is brittle and has shown to be worse at model sizes above $100\mathrm{M}$ parameters Tian et al. (2022). the applicability of model pre-training is currently limited due to the significantly lower model quality than in centralized training.
@@ 3.2 Parameter Efficient Fine-Tuning
Generally, PEFT is used to improve the performance of large models already trained on a large data corpus further and provide good performance across various tasks. This is especially effective since the data required for fine-tuning is orders of magnitude smaller than for pre-training. When applying PEFT, additional fully connected layers are inserted into the pre-trained model between the transformer blocks. While the original model weights are frozen, only the newly added layers are trained, typically resulting in $\geq 98\%$ less communication Hu et al. (2021). This renders PEFT techniques well-suited for FL applications since they address computation and communication alike. However, Babakniya et al. [2023] show in their study that PEFT is more sensitive towards non-IID data than FMT, but this sensitivity can be mitigated. PEFT can be applied in FL applications as follows.
Sparse fine-tuning of pre-trained model parameters.
As communication is a key concern in FL, reducing the number of parameters to communicate between client and server has become a priority. One of the most used approaches to achieve this is BitFit Zaken et al. (2021). The technique freezes almost all model parameters $w_{t}$ and only trains the bias term $b$ and the final classification layer $w^{\mathrm{final}}_{t}$ over the input features $a_{t}$, where the next-layer input features $a_{t+1}=a_{t}\cdot w_{t}+b$. With this technique, it is only required to communicate $b$ and $w^{\mathrm{final}}_{t}$. The communication load is reduced by $\geq 99\%$, i.e., instead of communicating 100 million parameters, only 100 thousand parameters are sent over the network.
@@
Adapter-based fine-tuning of additionally added parameters.
When aiming to maintain a pre-trained model while introducing task-specific knowledge, adapter-based fine-tuning techniques provide strong performance and on-par efficiency compared to sparse fine-tuning techniques Houlsby et al. (2019). Here are two adapter layers introduced in each transformer block of a foundation model. The adapter layer $a^{a}_{t+1}$ is calculated based on a downstream projection $w_{t}^{\mathrm{down}}$ of input feature $a^{a}_{t}$ into a lower-dimensional space $w_{t}^{\mathrm{down}}\in\mathcal{R}^{d\times r}$ followed by an upstream projection $w_{t}^{\mathrm{up}}\in\mathcal{R}^{r\times u}$, resulting in
$a^{a}_{t+1}=w_{t}^{\mathrm{up}}\cdot h(w_{t}^{\mathrm{down}}\cdot a^{a}_{t})\mathrm{.}$ (4)
With this, we can fine-tune an FM over much fewer dimensions than when fully fine-tuning a model. This saves both computational resources and communication costs in the same range as FedPEFT.
An improvement with regard to computational and communication efficiency over additional adapters is low-rank adapters (LoRA) Hu et al. (2021). The technique uses a lower dimensional representation $\mathbf{A}\in\mathcal{R}^{r\times u}$, where $u$ is the dimension of the next layer after the LoRA adapter and $\mathbf{B}\in\mathcal{R}^{d\times r}$, where $d$ is the dimension of the previous LoRA adapter. The weight updates are calculated with $w_{t+1}=w_{t}+\Delta w=w_{t}+\mathbf{B}\mathbf{A}$. $r\ll\min(d,u)$ casts the weight update matrices into a much lower dimensionality than in the original transformer module without the necessity of adding additional adapters, i.e., LoRA builds an adapter for existing parameters. However, as $\mathbf{A}$ is initialized randomly to a Gaussian distribution and $\mathbf{B}$ as a zero matrix, this works well for centralized settings with IID data Hu et al. (2021). For FL settings, this initialization method bears the risk of slowing down the fine-tuning process over non-IID data.
@@
Zhang et al. [2023b] provide a systematic benchmark study on adapter-based fine-tuning methods in privacy-preserving FL systems. Their results show that fine-tuning with additional adapters and LoRA both yield the same benchmark results regarding model accuracy. However, LoRA requires $66\%$ less communication than additional adapters.
However, both FedCLIP and FedPETuning yield a worse accuracy than full fine-tuning. SLoRA Babakniya et al. (2023) and FedDPA Yang et al. (2024) are LoRA-based techniques to fine-tune models in non-IID FL settings. Their approach parameterizes the weight update based on $r$,
@@
As $\beta$ depends on $r$, scaling the ratio helps control the weight update impact of a single client. Subsequently, this can be used to control inconsistent training updates caused by non-IID data. To practically achieve this, Babakniya et al. [2023] make use of a two-stage process: First, they used singular vector decomposition on $\mathbf{A}$ and $\mathbf{B}$ to obtain a common initialization point for LoRA across all clients in an FL system. Second, the training is facilitated with the commonly initialized low-rank representations. Their approach achieves on-par performance with full model fine-tuning. However, they require a warmup time of approximately 100 FL rounds for stage 1, which can be very expensive in FL settings as clients are often unavailable consecutively for such a large number of FL rounds McMahan et al. (2017).